The standard mass of an unknown substance is 15.6g. Which of the following set has the least accurate value? Show solutions



Trial A B C


1 15.6 15.8 15.6


2 15.4 14.6 15.5


3 15.6 14.7 15.8


1
Expert's answer
2021-09-27T09:03:17-0400

Find the mean of all these sets:

1) The mean is


"\mu=\frac{m_1+m_2+m_2}{3}=\frac{15.6+15.8+15.6}{3}=15.67\text{ g}."

2) The mean is


"\mu=\frac{m_1+m_2+m_2}{3}=\frac{15.4+14.6+15.5}{3}=15.17\text{ g}."

3) The mean is


"\mu=\frac{m_1+m_2+m_2}{3}=\frac{15.6+14.7+15.8}{3}=15.37\text{ g}."

The least accurate value is 2.